OpenStudy (anonymous):

Choose the correct simplification of the expression (8x2)3. 24x6 512x6 512x5 24x5

OpenStudy (anonymous):

is it in a power of 3 or multiply by 3..?

OpenStudy (anonymous):

the 2 and the 3 are powers.

jimthompson5910 (jim_thompson5910):

\[(8x^2)^3\] \[8^3x^{2\cdot 3}\] \[512x^6\] So \[(8x^2)^3=512x^6\]

OpenStudy (anonymous):

=\[=8^3 *x ^{(2*3)}=512 * x ^{6}\]